the volume of a cube is given by V=s^3, where s is the length of a side. find the length of a side of a cube if the volume is 500 ml^3. round the answer to 3 decimal places.
Answer by checkley77(12844)

V=s^3
500=s^3
c^3=500
s=cubert500
s=7.937 ans.